OpenStudy (anonymous):

someone help please HELP Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (wolfe8):

Well you know apple juice = 75% so water is 100-75 = 25% Keep in mind that percentage is the number over 100. So when you have 75% of apple juice, it means you have 75/100 of apple juice from the total mixture of apple juice and water. The same applies for 25% water. Have x(or anything you want) represent the total volume of mixture. So you have for apple juice: 75/100 * x water: 25/100 * x You know the water is 1 liter. So you have 25/100 * x = 1 This is how you set up the equations. Can you do it from here?

OpenStudy (anonymous):

I'm still kind of lost.. /:

OpenStudy (wolfe8):

What do you know about percentage?

OpenStudy (anonymous):

I know about percentage

OpenStudy (wolfe8):

Alright well maybe I can hint you: When I say 20% of 30 people, it means 20/100 *30 = 6 So 75% of a jar is apple juice means: 75/100 * jar(volume) You don't know what the total volume is so you will have to use a variable. That is what the question wants. I would say this: let x be the total volume of the jar. Then I can rewrite my apple juice equation as 75/100 * x Do the same for the water part.

OpenStudy (anonymous):

thank you !!